# A 4 - 8 GHz Kinetic Inductance Travelling-Wave Parametric Amplifier Using Four-Wave Mixing with Near Quantum-Limit Noise Performance

### 19 Feb 2024 v1

@@ I Introduction
Using a quantum noise-limited parametric amplifier as a first-stage amplifier for readout can improve the sensitivity of multiple cryogenic detector technologies, such as the Microwave Kinetic Inductance Detector (MKID) Zobrist et al. (2019), , and microwave SQUID multiplexer ($\mathrm{\mu}$MUX) readout of Mettalic Magnetic Calorimeters Kempf et al. (2017) and Transition-Edge Sensors Malnou et al. (2023a). Traveling wave parametric amplifiers are also an appealing choice for fast and high-fidelity readout Peng et al. (2022); Barzanjeh, DiVincenzo, and Terhal (2014); Didier et al. (2015) of cryogenic qubits. A commonly used frequency band to read out superconducting qubits and detectors is 4 - 8 GHz due to its availability, accessibility, and maturity of readout electronic systems and components at a relatively low cost. Applications of superconducting parametric amplifiers also extend to fundamental physics research for laboratory-based experiments such as dark matter searches, where a quantum-noise-limited gain is of interest Ramanathan et al. (2023).
@@
Kinetic inductance-based superconducting parametric amplifiers utilize the nonlinear kinetic inductance of superconducting thin films such as NbTiN. Four-wave mixing (4WM) processes can occur in kinetic inductance transmission line structures patterned in coplanar waveguide Eom et al. (2012) or microstrip line Shu et al. (2021) geometries when a strong pump tone is present. To obtain maximum gain over a desired bandwidth, a phase-matching condition is necessary. Implementing geometric dispersion allows for control of phase matching. This technique, called dispersion engineering, can also be utilized to mismatch unwanted higher-order nonlinear processes such as pump third harmonicEom et al. (2012).
@@
This paper presents a four-wave mixing KI-TWPA, where the pump tone is between the signal and idler frequencies, $2f_{pump}=f_{signal}+f_{idler}$. The amplifier tends to produce a gain in disjointed frequency ranges that we will refer to as the signal and idler (or image) bands, which makes it possible to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies. The devices described here are made of NbTiN and patterned in an inverted microstrip geometryShu et al. (2021). The transmission line is dispersion engineered by adding sinusoidally modulated capacitive stubs with a median length of 9.2 $\mathrm{\mu m}$ and modulation of 2.08 $\mathrm{\mu m}$ to create a bandgap centered at 12.5 GHz.
@@ II Device Design and Fabrication
$\Delta\beta=\kappa_{s}+\kappa_{i}-2\kappa_{p}=-\kappa_{p}\;\frac{I_{p}^{2}}{4I_{*}^{2}}$ (2)
$\kappa_{s}$,$\kappa_{i}$, and $\kappa_{p}$ are the wave vectors corresponding to the signal, idler, and pump tones, respectivelyEom et al. (2012).Fig. 1.c shows a plot of $\Delta\beta$ as a function of frequency.
Figure 1: a) A SEM micrograph of the transmission line with capacitive stubs. The width of the stubs is 340 nm, the spacing between them is 2.2 microns, and the average length of the stubs is 10.8 microns. b) Plot of the calculated propagation constant with a linear part subtracted as a function of frequency with a discontinuity at around 12.5 GHz, which results from the modulation of the length of the capacitive stubs with a periodicity of 122.7 microns. The inset shows the structure of the simulated dispersion around the bandgap and the position of the pump tone relative to the bandgap. c) Plot of calculated phase matching criterion, $\Delta\beta$, as a function of frequency. Shown in red is the self-phase modulation of the pump, and in purple is the difference between the pump wave number relative to the signal and idler wave numbers. The optimal parametric gain (maximum gain) is achieved near the outer intersecting points
@@
Figure 2: a) Shows the measured gain of the device at 1 K with pump frequency at 10.6 GHz. The frequencies over which the device produces gain are split into three bands. The gap at around 12 GHz is due to the “engineered” band gap produced by the periodic modulation of the transmission line structureShu et al. (2021); Klimovich (2022). The gap around 8 GHz corresponds to frequencies where the idler tone ends up in the band gap, so the 4WM process is inoperative. The dashed line shows the simulated gained curve with the same pump condition as the measured gain of the device. b) plot of measured gain of the device at 1 K with different pump frequencies. The bandwidth of the “signal” band can be adjusted significantly by varying the pump frequency.
@@ III.1 Behavior
The gain is measured by normalizing the pumped transmission by the pump-off transmission. We expect negligible loss from the KI-TWPA based on results reported in Shu et al.Shu et al. (2021) for a similar device geometry in this frequency range. The KI-TWPA tends to produce gain in disjointed frequency ranges that are referred to as the signal and idler bands (see Fig. 1.a). The amplifier also produces some gain in a narrow frequency band around the pump frequency, but the gain there is less due to a lack of phase matching. The separation of the bands has the advantage that it is easier to use a diplexer to separate out the idler band.The expected gain of the device was calculated by solving the coupled-mode equations for a degenerate four-wave mixing process Shu et al. (2021). Comparing the simulation plot shown in Fig. 2.a indicates a good agreement with measurement results plotted in the same figure. The peak of the signal gain of the device can be tuned by adjusting the pump tone frequency at different values below the bandgap frequency as governed by the following relation
$\omega_{s}=2\omega_{p}-\omega_{i}.$ (3)
$\omega_{s}$, $\omega_{p}$, and $\omega_{i}$ are the signal, pump, and idler frequencies, respectively.
@@
To measure the compression point, a test tone in the middle of the device’s signal band was injected into the KI-TWPA, and the system’s output power was measured using a spectrum analyzer. With the pump frequency set to 10.6 GHz and pump tone power of -23 dBm, resulting in 15 dB of gain, the output power of the amplified test tone at the spectrum analyzer was measured as a function of the test tone’s power at the device. As shown in Fig. 3.a, the gain compresses by 1 dB when the input signal at the device is $\simeq$ -58 dBm.
@@
Figure 3: a) Plot of the measured output power at the spectrum analyzer of a test tone as a function of the test tone’s input power at 20 mK with a signal gain of 15 dB. The gain compresses by 1 dB from the linear response when the input signal power reaches -58 dBm. b) Plot of the measured gain for a fixed pump tone frequency as a function of pump tone power at 20 mK. Both of these measurements were taken using the setup shown in Fig. 5.
@@
Figure 4: a) Schematic of the noise measurement setup with hot and cold broadband noise sources. Two 50 $\mathrm{\Omega}$ cryogenic loads were used as a noise source. The hot source was heat sunk to the 3 K stage of the dilution fridge with a base temperature of 3.18 K, and the cold noise source was placed on the fridge’s mixing chamber (MXC) with a base temperature of $\simeq 20$ mK. All other components in this figure were placed in the MXC. Using a relay switch S, the input of the KI-TWPA was switched between the hot and cold sources. The two types of diplexers have crossovers in the lower and upper gaps in the gain curve shown in Fig. 2.a and are used to inject and separate out the pump and idler tones. c) circuit diagram of the cascaded noise model for the KI-TWPA measurement setup.
@@ III.2 Noise
The amplifier chain noise was measured using a Y-factor method with two 50 $\mathrm{\Omega}$ cryogenic loads as noise sources, one at 3.18 K and the other at the Mixing Chamber (MXC) stage ($\sim$ 20 mK) of a dilution refrigerator. Using a relay switch, we alternate between the noise sources as shown in Fig. 4.a. Due to a lack of output-to-input isolation, reflections must be avoided at either side of the TWPA at frequencies where the amplifier has gain. If the sum of the dB return losses on the input and output sides does not exceed the gain, an oscillation will develop. The noise measurement setup explained here achieves this using a combination of diplexer (DPX) circuits. The input side diplexers allow for the injection of the pump tone through a bandpass filter that removes synthesizer phase noise. The input side DPX I provides the input cold idler frequency termination ($>$13 GHz). The noise from that termination determines the added noise from the KI-TWPA, so it is kept at a temperature $T\ll\hbar\omega_{i}/k_{B}$ for the system to be quantum limited. This condition is met by terminating the idler frequencies at the MXC ($\sim 20$ mK).In this configuration, the KI-TWPA only receives the broadband white noise emitted from the sources in the signal band.
On the output side, the diplexers remove the pump and idler tones and terminate them. The pump is terminated at a higher temperature stage ($\sim$ 50 mK) to minimize heating. The idler is terminated at the base temperature to avoid noise from that termination, which may travel back through the KI-TWPA from output to input and possibly be reflected back into the KI-TWPA. Finally, an isolator is used to avoid noise from the HEMT, which may otherwise contaminate the KI-TWPA via the same mechanism. the noise spectrum is then read through a room-temperature low noise amplifier (LNA) and measured using a spectrum analyzer.
The total system noise spectrum in units of quanta, which includes the KI-TWPA, a Low Noise Factory cryogenic HEMT amplifier (Model: LNF-LNC0.3_14B), and a room temperature low noise amplifier, was measured using the Y-factor method,
$N_{sys}(\omega)=\frac{T_{H}(\omega)-Y\;T_{C}(\omega)}{Y-1}\Bigg{(}\frac{k_{B}}{\hbar\;\omega}\Bigg{)},$ (4)
where $\omega$ is the signal angular frequency, $k_{B}$ is the Boltzmann constant, and $\hbar$ is the reduced Planck’s constant. The $T_{C}(\omega)$ and $T_{H}(\omega)$ are the effective temperatures of the hot and cold terminations given by Kerr and Feldman (1996)
$T_{H,C}=\frac{\hbar\omega}{2k_{B}}\coth\Big{(}\frac{\hbar\omega}{K_{B}T^{\prime}_{H,C}}\Big{)}$ (5)
@@
The KI-TWPA added noise was measured using the gain of the KI-TWPA and the total system noise. However, microwave losses in the measurement setup, especially between the KI-TWPA and the noise sources, can cause an overestimation of the measured noise. We estimated the loss from the diplexers by comparing the measured transmission through the KI-TWPA and diplexers to a cable bypassing them. After considering the loss, the added noise of the KI-TWPA can be calculated using the following equation.
@@
where $N_{A}$ is the added noise of the KI-TWPA, $L_{1}$ and $L_{2}$ are the losses in components before and after the KI-TWPA, respectively. $N_{qm}$ is the quantum noise, $N_{HEMT}$ is the added noise of the HEMT amplifier and $G_{PA}$ is the measured gain of the KI-TWPA. The noise model used to arrive at equation (6) is explained in Appendix B. We measure a total system noise of $1\leq N_{sys}\leq 2$ in the 4.6 - 8.1 GHz range. Using equation (6), we estimate the added noise of the KI-TWPA to be $0.5\leq N_{sys}\leq 1.5$ as shown in Fig. 4. We can also estimate the cryogenic HEMT amplifier’s noise from unpumped KI-TWPA measurements, where we measure an average noise temperature of 5 K or $\sim$ 13 quanta in the 4 - 8 GHZ band consistent with data provided in the amplifier’s datasheet. It was observed that fluctuations in the measured noise spectra and added noise in TWPA were more than half a quanta. These fluctuations could be attributed to some nonidealities in the test setup that could be improved. For instance, both noise sources are connected to the input of the KI-TWPA through three cables and two switches. Estimating the insertion loss in all these components is quite challenging, and any unaccounted losses can lead to uncertainty in the measurement and overestimation of the measured noise. To address this issue, we can use a Variable Temperature Source (VTS) or a shut-noise tunnel junction (SNTJ) Malnou et al. (2023b). Using a VTS would require only one 50 $\Omega$ termination and could be physically placed closer to the KI-TWPA. This could lead to fewer cables required between the noise source and KI-TWPA, which would improve the uncertainty and loss of the noise power at the input of the KI-TWPA.
@@ IV Conclusion
In conclusion, we have designed, fabricated, and measured a wide-bandwidth Four-Wave Mixing KI-TWPA that produces gain between 3 - 9 GHz with a 1 dB compression of -58 dBm suitable for the readout of large detector arrays and superconducting qubits. We have demonstrated a near quantum-limited noise performance of KI-TWPA in the 4.5 to 8 GHz frequency range using a y-factor method.
The signal band peak gain of this device is shown to be tunable with the frequency of the pump tone. This tunability makes the KI-TWPA very robust for applications where the signal gain needs to be adjusted, such as in dark matter experiments. Moving forward, KI-TWPA performance could benefit from using lower $T_{c}$ materials with high kinetic inductance. For instance, a KI-TWPA with a TiN or WSi signal layer would require lower pump power to achieve parametric gain. We also intend to design and integrate superconducting diplexer circuits with the KI-TWPA for compactness and robustness.

### 13 Mar 2024 v3

@@ I Introduction
Using a quantum noise-limited parametric amplifier as a first-stage amplifier for readout can improve the sensitivity of multiple cryogenic detector technologies, such as the Microwave Kinetic Inductance Detector (MKID) Zobrist et al. (2019), , and microwave SQUID multiplexer ($\mathrm{\mu}$MUX) readout of Mettalic Magnetic Calorimeters Kempf et al. (2017) and Transition-Edge Sensors Malnou et al. (2023). Traveling wave parametric amplifiers are also an appealing choice for fast and high-fidelity readout Peng et al. (2022); Barzanjeh, DiVincenzo, and Terhal (2014); Didier et al. (2015) of cryogenic qubits. A commonly used frequency band to read out superconducting qubits and detectors is 4 - 8 GHz due to its availability, accessibility, and maturity of readout electronic systems and components at a relatively low cost. Applications of superconducting parametric amplifiers also extend to fundamental physics research for laboratory-based experiments such as dark matter searches, where a quantum-noise-limited gain is of interest Ramanathan et al. (2023).
@@
Kinetic inductance-based superconducting parametric amplifiers utilize the nonlinear kinetic inductance of superconducting thin films such as NbTiN. Four-wave mixing (4WM) processes can occur in kinetic inductance transmission line structures patterned in coplanar waveguide Eom et al. (2012) or microstrip line Shu et al. (2021) geometries when a strong pump tone is present. A phase-matching condition is necessary to obtain maximum gain over a desired bandwidth. Implementing geometric dispersion allows for control of phase matching. This technique, called dispersion engineering, can also be utilized to mismatch unwanted higher-order nonlinear processes such as pump third harmonicEom et al. (2012).
@@
This paper presents a degenerate four-wave mixing KI-TWPA, where the pump tone is between the signal and idler frequencies, $2f_{pump}=f_{signal}+f_{idler}$. The amplifier tends to produce a gain in disjointed frequency ranges that we will refer to as the signal and idler (or image) bands, which makes it possible to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies. The devices described here are made of NbTiN and patterned in an inverted microstrip geometryShu et al. (2021). The transmission line is dispersion engineered by adding sinusoidally modulated capacitive stubs with a median length of 9.2 $\mathrm{\mu m}$ and modulation of 2.08 $\mathrm{\mu m}$ to create a bandgap centered at 12.5 GHz.
@@ II Device Design and Fabrication
$\Delta\beta=\kappa_{s}+\kappa_{i}-2\kappa_{p}=-\kappa_{p}\;\frac{I_{p}^{2}}{4I_{*}^{2}}\,,$ (2)
where $\kappa_{s}$,$\kappa_{i}$, and $\kappa_{p}$ are the wave vectors corresponding to the signal, idler, and pump tones, respectively, and the term on the right represents the nonlinear dispersive effect Eom et al. (2012).Fig. 1.c shows a plot of $\Delta\beta$ as a function of frequency.
Figure 1: a) A SEM micrograph of the transmission line with capacitive stubs. The width of the stubs is 340 nm, the spacing between them is 2.2 microns, and the average length of the stubs is 10.8 microns. b) Plot of the calculated propagation constant with a linear part subtracted as a function of frequency with a discontinuity at around 12.5 GHz, which results from the modulation of the length of the capacitive stubs with a periodicity of 122.7 microns. The inset shows the structure of the simulated dispersion around the bandgap and the position of the pump tone relative to the bandgap. c) Plot of calculated phase matching criterion, $\Delta\beta$, as a function of frequency. Shown in red is the self-phase modulation of the pump,$\kappa_{p}\frac{I_{p}^{2}}{4I_{*}^{2}}$, and in purple is the difference between the pump wave number, $2\kappa_{p}$, relative to the signal and idler wave numbers, $\kappa_{s}$ and $\kappa_{i}$. The optimal parametric gain (maximum gain) is achieved near the outer intersecting points.
@@
Figure 2: a) Shows the measured gain of the device at 1 K with pump frequency at 10.6 GHz. The frequencies over which the device produces gain are split into three bands. The gap at around 12 GHz is due to the “engineered” band gap produced by the periodic modulation of the transmission line structureShu et al. (2021); Klimovich (2022). The gap around 8 GHz corresponds to frequencies where the idler tone ends up in the band gap, so the 4WM process is inoperative. The dashed line shows the simulated gained curve with the same pump condition as the measured gain of the device. b) plot of measured gain of the device at 1 K with different pump frequencies. The bandwidth of the “signal” band can be adjusted significantly by varying the pump frequency. Measurements shown in this sub-figure are from another device on the same wafer with an identical design.
@@ III.1 Behavior
The gain is measured by normalizing the pumped transmission by the pump-off transmission. We expect negligible loss from the KI-TWPA based on results reported in Shu et al.Shu et al. (2021) for a similar device geometry in this frequency range. The KI-TWPA tends to produce gain in disjointed frequency ranges that we refer to as the signal and idler bands (see Fig. 1.a). The amplifier also produces some gain in a narrow frequency band around the pump frequency, but the gain there is less because, in that frequency range, the nonlinear dispersive effect is not compensated by the engineered dispersion. The gap in the gain curve around 12.5 GHz is simply the result of the transmission gap due to the modulation of the stub lengths. The additional gap around 8.7 GHz occurs because with the signal tone near that frequency, the idler tone would fall within the bandgap, and therefore, the gain process is inoperative. The gaps in the gain curve serve a useful function in that it is easier to use a diplexer to separate out the pump and the signal and idler bands. Typically, diplexers have poor input matching at their crossover frequencies, so by placing those frequencies in the zero gain regions, potential problems with reflections can be avoided.
The expected gain of the device was calculated using the low power dispersion extracted from the measured unpumped transmission. The measurement was compared to a network model of the KI-TWPA to determine the film surface inductance. In our method, described in ref. Shu et al. (2021), the modulated transmission line is considered to be a uniform medium with an effective dielectric constant, which we derive from the adjusted network model. That effective dielectric constant is used in the coupled-mode equations to predict the gain. For the simulated gain curve shown in Fig. 2.a, the higher-frequency products at $2\omega_{p}+\omega_{s}$, $3\omega_{p}$, and $4\omega_{p}-\omega_{s}$ were included in the calculation in addition to the signal, pump and idler frequenciesShu et al. (2021). To achieve a similar gain level to the measured value, we adjusted the value of the pump current used in the calculation. Comparing the simulation plot shown in Fig. 2.a indicates a good agreement with measurement results.
The frequency band over which the device produces gain can be tuned significantly by adjusting the frequency of the pump tone, as shown in Fig. 2.b. As the dispersion introduced by the periodic modulation changes rapidly below the bandgap (Fig. 1.b), small pump frequency changes result in somewhat larger changes in the signal and idler frequencies that satisfy the phase matching condition, eqn. 2.
@@
To measure gain compression, a test tone in the middle of the device’s signal band was injected into the KI-TWPA, and the system’s output power was measured using a spectrum analyzer. With the pump frequency set to 10.6 GHz and pump tone power of -23 dBm, resulting in 15 dB of gain, the output power of the amplified test tone at the spectrum analyzer was measured as a function of the test tone’s power at the device. As shown in Fig. 3.a, the gain compresses by 1 dB when the input signal is $\simeq$ -58 dBm, which corresponds to an output power of -43 dBm. As the power of the output signal and idler tones is provided by the pump, compression occurs when the pump is level significantly depleted. For our device, that occurs with an output signal power of about -20 dB below the pump power.
@@
Figure 3: a) Plot of the measured output power at the spectrum analyzer of a test tone as a function of the test tone’s input power at 20 mK with a signal gain of 15 dB. The gain compresses by 1 dB from the linear response when the input signal power reaches -58 dBm. b) Plot of the measured gain for a fixed pump tone frequency as a function of pump tone power at 20 mK. Both of these measurements were taken using the setup shown in Fig. 5. The higher ripple level in this plot compared to Fig. 2.a is due to the addition of more components in the measurement setup, such as the diplexers and more cabling, which can cause impedance mismatch.
@@
Figure 4: a) Schematic of the noise measurement setup with hot and cold broadband noise sources. Two 50 $\mathrm{\Omega}$ cryogenic loads were used as noise sources. The hot source was heat sunk to the 3 K stage of the dilution fridge with a base temperature of 3.18 K, and the cold noise source was placed on the fridge’s mixing chamber (MXC) with a base temperature of $\simeq 20$ mK. All other components in this figure were placed in the MXC. Using a relay switch S, the input of the KI-TWPA was switched between the hot and cold sources. The two types of diplexers have crossovers in the lower and upper gaps in the gain curve shown in Fig. 2.a and are used to inject and separate out the pump and idler tones. c) circuit diagram of the cascaded noise model for the KI-TWPA measurement setup.
@@ III.2 Noise
The amplifier chain noise was measured using a Y-factor method with two 50 $\mathrm{\Omega}$ cryogenic loads as noise sources, one at 3.18 K and the other at the Mixing Chamber (MXC) stage ($\sim$ 20 mK) of the dilution refrigerator. Using a cryogenic relay switch, we alternate between the noise sources as shown in Fig. 4.a. Due to a lack of output-to-input isolation, reflections must be avoided at either side of the TWPA at frequencies where the amplifier has gain. If the sum of the dB return losses on the input and output sides does not exceed the gain, an oscillation will develop. The circuit shown in Fig. 4 provides an impedance match at both the input and output of the KI-TWPA over the signal and idler frequency ranges using a combination of diplexer (DPX) circuits. The input side diplexers allow for the injection of the pump tone through a bandpass filter (not shown) that removes synthesizer phase noise. The input side DPX I provides a cold termination to the input of the KI-TWPA over the idler band ($>$13 GHz). The noise from that termination determines the added noise of the KI-TWPA, so it is kept at a temperature $T\sim 20\,$mK $\ll\hbar\omega_{i}/k_{B}$ for the system to be quantum limited. In this configuration, the KI-TWPA only receives the thermal noise emitted from the calibration sources in the signal band.
On the output side of the KI-TWPA, the diplexer removes the pump and idler tones and terminates them. To minimize heating, the pump is terminated at a higher temperature stage ($\sim$ 50 mK). The idler is also terminated at 50 mK to keep the KI-TWPA quantum limited. Finally, an isolator is used to avoid noise from the HEMT, which may otherwise contaminate the KI-TWPA via the same mechanism. The signal is then further amplified using a room temperature amplifier and measured using a spectrum analyzer.
The total noise of the amplifier chain, which includes the KI-TWPA, a Low Noise Factory cryogenic HEMT amplifier (Model: LNF-LNC0.3_14B), and a room temperature low noise amplifier, was determined using the Y-factor method,
$N_{sys}(\omega)=\frac{N_{H}(\omega)-Y\;N_{C}(\omega)}{Y-1},$ (3)
where $\omega$ is the signal angular frequency, $k_{B}$ is the Boltzmann constant, and $\hbar$ is the reduced Planck’s constant. The $N_{C}(\omega)$ and $N_{H}(\omega)$ are the noise temperatures of the hot and cold terminations in units of quanta given by Kerr and Feldman (1996)
$N_{H,C}=\frac{1}{2}\coth\Big{(}\frac{\hbar\omega}{2k_{B}T^{\prime}_{H,C}}\Big{)}$ (4)
@@
The noise added by the KI-TWPA itself was determined using the measured gain of the device and the total system noise from the Y-factor measurement. Microwave losses in the measurement setup, especially between the KI-TWPA and the noise sources, can cause an overestimation of the measured noise. We estimated the loss from the diplexers by comparing the measured transmission through the KI-TWPA and diplexers to that of a bypass cable as shown in Fig. 5.a. After considering the loss, the added noise, $N_{A}$, in units of quanta, of the KI-TWPA can be calculated as
@@
where $L_{1}$ and $L_{2}$ are the losses in the components before and after the KI-TWPA, respectively. $N_{qm}=0.5$ is the quantum noise, $N_{HEMT}$ is the added noise of the HEMT amplifier and $G_{PA}$ is the measured gain of the KI-TWPA. The noise model used to arrive at equation (6) is explained in Appendix B. We measure a total system noise of $1\leq N_{sys}\leq 2$ in the 4.6 - 8.1 GHz range as shown in Fig. 4.c. Using equation (6), we estimate the added noise of the KI-TWPA to be $0.5\leq N_{sys}\leq 1.5$ as shown in Fig. 4. The cryogenic HEMT amplifier’s noise was measured using the Y-factor with the pump off. We measured an average noise temperature of 5 K or $\sim$ 13 quanta in the 4 - 8 GHZ band consistent with data provided in the amplifier’s datasheet. It was observed that the measured added noise in the KI-TWPA deviates from half a quanta in some places (Fig. 4.c). These discrepancies could be attributed to some nonidealities in the test setup. For instance, both noise sources are connected to the input of the KI-TWPA through three cables and two switches. Estimating the insertion loss in all these components is quite challenging, and any unaccounted losses can lead to uncertainty in the measurement and overestimation of the measured noise. Impedance mismatch between the noise terminations and the KI-TWPA is also another source of uncertainty in our noise measurements. The ripples in the measured system noise and the TWPA-added noise are indications of these imperfections.
@@ IV Conclusion
In conclusion, we have designed, fabricated, and measured a wide-bandwidth Four-Wave Mixing KI-TWPA that produces gain between 3 - 9 GHz with a 1 dB compression of -58 dBm suitable for the readout of large detector arrays and superconducting qubits. Using a Y-factor method, we have demonstrated a near quantum-limited noise performance of KI-TWPA in the 4.6 to 8.1 GHz frequency range. The amplifier produces gain in disjointed frequency ranges that we call the signal band and a higher-frequency idler (or image) band. This makes it straightforward to separate the idler using ancillary circuitry and use the full signal band without contamination from the image (or idler) frequencies for the readout of cryogenic detectors. The signal band peak gain of this device is shown to be tunable with the frequency of the pump tone. This tunability makes the KI-TWPA suitable for applications where the signal frequency range needs to be adjusted, such as in dark matter experiments. In addition, The higher frequency idler band of the KI-TWPA can be used to look for dark matter candidates with a higher mass range, which makes it possible to cover a wide range of masses with a single KI-TWPA.